OpenStudy (anonymous):

what number is less twice its reciprocal gives 23/5?

OpenStudy (anonymous):

x-(2/x)=23/5 multiply everything by x x^2-2=(23/5)x multiply everything by 5 5x^2-10=23x rearrange 5x-23x-10=0 solve using the quadratic formula x=5 5-(2/5)=23/5